# Instruct-MusicGen: Unlocking Text-to-Music Editing for Music Language Models via Instruction Tuning

### v2

@@ Abstract
Recent advances in text-to-music editing, which employ text queries to modify music (e.g. by changing its style or adjusting instrumental components), present unique challenges and opportunities for AI-assisted music creation. Previous approaches in this domain have been constrained by the necessity to train specific editing models from scratch, which is both resource-intensive and inefficient; other research uses large language models to predict edited music, resulting in imprecise audio reconstruction. To Combine the strengths and address these limitations, we introduce Instruct-MusicGen, a novel approach that finetunes a pretrained MusicGen model to efficiently follow editing instructions such as adding, removing, or separating stems. Our approach involves a modification of the original MusicGen architecture by incorporating a text fusion module and an audio fusion module, which allow the model to process instruction texts and audio inputs concurrently and yield the desired edited music. Remarkably, Instruct-MusicGen only introduces $\sim$8new parameters to the original MusicGen model and only trains for 5K steps, yet it achieves superior performance across all tasks compared to existing baselines, and demonstrates performance comparable to the models trained for specific tasks. This advancement not only enhances the efficiency of text-to-music editing but also broadens the applicability of music language models in dynamic music production environments.  [^1] Work done during Yixiao’s internship at Sony AI. [^2] Code, model weights and demo are available at: https://bit.ly/instruct-musicgen.
@@ 1 Introduction
The rapid advances in text-to-music generation have opened up new possibilities for AI-assisted music creation (Agostinelli et al., 2023; Copet et al., 2023; Li et al., 2023; Liu et al., 2023a; Chen et al., 2023). This paradigm shift has also sparked a growing interest in developing models that offer greater controllability (Lin et al., 2023; Wu et al., 2023a; Melechovsky et al., 2023; Lin et al., 2024) and editability (Han et al., 2023; Zhang et al., 2024; Hussain et al., 2023) over the music generation process. In music production, a stem—a mixed group of tracks often related by instrument type (like drums or lead vocals)—is essential for mixing and mastering because it allows producers to isolate, adjust, and manipulate individual elements of a song. Following the definition in  Zhang et al. (2024), "text-to-music editing" involves using textual queries to modify various aspects of a music recording, which can be categorised into two main types: intra-stem editing, which focuses on modifying a single stem (e.g., changing the instrument, timbre, or performance style), and inter-stem editing, which involves altering the relationships among stems (e.g., adding, removing, or separating stems). Our work mainly focuses on the problem of inter-stem editing.
Previous attempts to develop text-based music editing models have encountered several challenges. Some approaches (Wang et al., 2023; Han et al., 2023) have focused on training specialized editing models from scratch, which is resource-intensive and may not yield results comparable to state-of-the-art music generation models. Other work (Zhang et al., 2023b; Hussain et al., 2023; Yang et al., 2023) has sought to leverage existing language models (LLMs) and MusicGen (Copet et al., 2023), allowing the LLM to interpret editing instructions without further training the music model. Although this approach offers flexibility, it often lacks the ability to precisely reconstruct the conditional audio, leading to unreliable results. To address these limitations, an ideal solution should harness the knowledge embedded in pretrained models to ensure high-quality audio output while adapting the architecture to accommodate the specific requirements of music editing tasks.
@@
In this paper, we introduce Instruct-MusicGen, a novel approach that applies an instruction-following tuning strategy to the pretrained MusicGen model, enhancing its ability to follow editing instructions effectively without finetuning all its parameters. As shown in Figure 1, by incorporating an audio fusion module based on LLaMA-adapter (Zhang et al., 2023a; Lin et al., 2023) and a text fusion module based on LoRA (Hu et al., 2021) into the original MusicGen architecture, we allow the model to process both precise audio conditions and text-based instructions simultaneously, while the original MusicGen does not. This enables Instruct-MusicGen to perform a wide range of editing tasks, such as adding, separating, and removing stems. To train Instruct-MusicGen, we synthesize an instructional dataset using the Slakh2100 dataset (Manilow et al., 2019), introducing only  8% additional parameters compared to the original model, and finetune the model for only 5K steps, which is less than 1% of training a music editing model from scratch.
We evaluate Instruct-MusicGen on two datasets: the Slakh test set and the out-of-domain MoisesDB dataset (Pereira et al., 2023). Our model outperforms existing baselines and achieves performance comparable to models specifically trained for individual tasks. This demonstrates the effectiveness of our approach in leveraging pretrained models for text-to-music editing while maintaining high-quality results. The main contributions of this work are threefold:
  1. 1.
    We propose Instruct-MusicGen, a novel approach that finetunes pretrained large music language models to follow editing instructions effectively. This method addresses the limitations of previous approaches by leveraging pretrained models to ensure high-quality audio output and precise reconstruction, combining both music and text multi-modal control.
  2. 2.
    We enhance the capabilities of a pretrained text-to-music model to perform multiple tasks, including adding, separating, and extracting stems from music audio, all within a single training process. This approach significantly reduces the computing resources cost compared to training specialised editing models from scratch.
  3. 3.
    We extend the benchmarking for text-to-music editing tasks, providing a comprehensive comparison across various models. Our model outperforms existing baselines and achieves performance comparable to models specifically trained for individual tasks, demonstrating the effectiveness and versatility of our approach.
@@ 2 Related work
Despite the growing body of research on text-to-music generation (Agostinelli et al., 2023; Copet et al., 2023; Li et al., 2023; Yao et al., 2023; Melechovsky et al., 2023; Evans et al., 2024a, b; Chen et al., 2023), the creation of music remains an inherently iterative process, involving multiple rounds of editing and revision. This ongoing need for refinement and adjustment underscores the necessity for specialized models that facilitate and enhance music editing.
Text-based music editing provides a flexible approach for editing music using textual queries. This method is similar to other modalities that require editing, such as image (Brooks et al., 2023; Liu et al., 2023b) and video (Chai et al., 2023; Ceylan et al., 2023) editing. In text-to-music editing, tasks leverage text to make precise alterations to existing music compositions. To address these needs in music editing, AUDIT (Wang et al., 2023) and InstructME (Han et al., 2023) developed a diffusion model trained with paired music editing data. Additionally, models like M2UGen Hussain et al. (2023), Loop Copilot (Zhang et al., 2023b), MusicAgent (Yu et al., 2023), ComposerX (Deng et al., 2024) and WavCraft (Liang et al., 2024) use large language models (LLMs) for reasoning and regenerate music with external music generation models. Furthermore, Postolache et al. (2024) attempt to jointly model multi-stem distribution of music for text-based generation and separation. Certain models focus exclusively on specific tasks within music editing, such as conditioned generation (Lin et al., 2023, 2024) and separation (Liu et al., 2023c), along with intra-stem editing tasks like text-based timbre transfer and style transfer (Zhang et al., 2024; Manor and Michaeli, 2024; Li et al., 2024).
The task of inter-stem music editing is closely related to stem-wise music generation, which is also crucial in the music production process. While these efforts are not directly tied to text-based controls, some research focuses on modeling stem-wise representations to enable simultaneous stem generation and separation. For instance, Jen-1 Composer (Yao et al., 2023) and MSDM (Mariani et al., 2023) jointly model the distribution of music with four stems using a diffusion model. The abilities of most existing stem-wise music models are restricted to a fixed set of 4 stems, which limits flexibility but enhances controllability; In contrast, StemGen (Parker et al., 2023) introduces a novel approach by employing a non-autoregressive Transformer architecture for more flexible stem-wise generation, moving beyond the constraints of a fixed stem set.
Our work distinguishes itself from these existing efforts in several key ways. First, rather than developing a new model from scratch or strictly adhering to a fixed set of stems, we leverage the power of a pretrained music language model, MusicGen, and enhance it with instruction tuning. This approach not only reduces the computational cost but also retains the high audio quality of the original MusicGen model. By integrating both audio and text fusion modules, our model, Instruct-MusicGen, can precisely interpret and execute complex editing instructions involving multiple stems. Furthermore, our method introduces minimal additional parameters and requires significantly less training, demonstrating a more efficient and scalable solution for text-based music editing. This dual approach of using pretrained models with specialized adapters sets our work apart, providing a flexible and powerful tool for the evolving needs of AI-assisted music creation.
@@ 3.1 MusicGen
The original MusicGen consists of three components: (1) the EnCodec (Défossez et al., 2022) audio encoder and decoder, which compress music audio waveforms into latent codes and reconstruct them back into waveforms; (2) a multi-layer transformer architecture that models sequences of latent codes, capturing higher-level music representations and efficiently modeling internal relationships within music audio; and (3) the T5 (Raffel et al., 2020) text encoder, which converts text descriptions into embeddings for text-conditioned generation.
MusicGen uses EnCodec (Défossez et al., 2022), a convolutional auto-encoder, to map real audio waveforms to latent codes. EnCodec employs Residual Vector Quantization (RVQ) (Zeghidour et al., 2022) to compress audio into tokens using multiple codebooks, where each quantizer encodes the quantization error from the previous one. For a reference audio $X\in\mathbb{R}^{d\cdot f_{n}}$, where $d$ is the duration and $f_{n}$ is the sample rate, EnCodec compresses $X$ into $Q\in\{1,...,L\}^{N\times d\cdot f_{s}}$, where $L$ is the RVQ codebook size, $N$ is the number of codebooks, and $f_{s}$ is the latent code sample rate ($f_{s}<<f_{n}$). In MusicGen, $N=4$, $f_{n}=50$, $f_{s}=32000$, and $L=2048$. Finally, The transformer models the sequence relationships over latent codes.
@@ 3.2 Instruct-MusicGen
Instruct-MusicGen takes a music audio input $X_{\text{cond}}$ and a text instruction $X_{\text{instruct}}$ (e.g., "Add guitar") as inputs. The model then edits the music audio $X_{\text{cond}}$ according to the instruction $X_{\text{instruct}}$ and generates the desired edited music $X_{\text{music}}$. As illustrated in Figure 2, Instruct-MusicGen incorporates two additional modules into the vanilla MusicGen: an audio fusion and a text fusion module.
Figure 2: The architecture of instruct-MusicGen. The audio fusion module transforms the conditional music audio into embeddings using a duplicated encoder and integrates these embeddings into the MusicGen decoder. The text fusion module modifies the cross-attention mechanism to handle text instructions by finetuning specific layers while keeping the text encoder parameters frozen.
@@ 3.2.1 Audio Fusion Module
The audio fusion module enables Instruct-MusicGen to accept external audio inputs, which is inspired by Zhang et al. (2023a) and Lin et al. (2023). Figure 2 presents the diagram of the audio fusion module. Initially, we convert $X_{\text{cond}}$ into EnCodec tokens, followed by re-encoding these tokens into the embedding $z_{\text{cond}},z_{\text{cond}}$ through the pre-trained embedding layers of MusicGen. Similarly, we transform $X_{\text{music}}$ into the pretrained embedding $z_{\text{music}}$.
@@
The module begins with duplicating self-attention modules of the pretrained MusicGen model to extract latent representations of $z_{\text{cond}}$. Given that MusicGen consists of $M$ layers, we denote $Z_{\text{cond}}=\{z^{0}_{\text{cond}},z_{\text{cond}}^{1},\ldots,z_{\text{cond}}^{M}\}$ and $Z_{\text{music}}=\{z^{0}_{\text{music}},z_{\text{music}}^{1},\ldots,z_{\text{music}}^{M}\}$ to represent the hidden states of $z_{\text{cond}}$ and $z_{\text{music}}$ respectively. Note that we use a learnable input embedding as $z^{0}_{\text{cond}}$ and initialize $z^{0}_{\text{music}}$ with $z_{\text{music}}$.
Let $W_{\rm Q}$, $W_{\rm K}$, $W_{\rm V}$ and $W_{\rm O}$ denote the query, key, value, and output projection matrices of the pretrained self-attention transformer layer respectively, then, $z_{\text{music}}$ is projected into the latent space,
$K=Z_{\text{music}}W_{\text{K}},\hskip 14.22636ptQ=Z_{\text{music}}W_{\text{Q}},\hskip 14.22636ptV=Z_{\text{music}}W_{\text{V}}.$ (1)
The attention output $O$ of $Q$, $K$, and $V$ is computed as follows,
$O={\rm Attention}(Q,K,V)W_{\rm O},$ (2)
${\rm Attention}(Q,K,V)=\text{Softmax}\left(\frac{QK^{\mathsf{T}}}{\sqrt{d}}\right)V,$ (3)
where $d$ is the dimension of $K$.
For $z_{\text{cond}}$, we design a learnable linear layer $\text{Linear}_{\text{cond}}$ to incorporate information from $X_{\text{cond}}$ into MusicGen, and $M$ layer-wise learnable position embeddings $\{e_{1},e_{2},....e_{M}\}$ to enhance the sequital modeling. Let $Z^{\prime}_{\text{cond}}=\{z_{\text{cond}}^{1}+\text{Linear}_{\text{cond}}(z_{\text{cond}})+e_{1},z_{\text{cond}}^{2}+\text{Linear}_{\text{cond}}(z_{\text{cond}})+e_{2},\ldots,z_{\text{cond}}^{M}+\text{Linear}_{\text{cond}}(z_{\text{cond}})+e_{M}\}$, then, $Z^{\prime}_{\text{cond}}$ is projected into the latent space,
$K^{\prime}=Z^{\prime}_{\text{cond}}W_{\text{K}},\hskip 14.22636ptQ^{\prime}=Z^{\prime}_{\text{cond}}W_{\text{Q}},\hskip 14.22636ptV^{\prime}=Z^{\prime}_{\text{cond}}W_{\text{V}}.$ (4)
We compute the attention output of $Q^{\prime}$, $K^{\prime}$, and $V^{\prime}$ and pass it through each layer sequentially. Let ${Q^{\prime}}_{m}$, ${K^{\prime}}_{m}$, and ${V^{\prime}}_{m}$ denote $m$-th layer value in $Q^{\prime}$, $K^{\prime}$, and $V^{\prime}$, which are computed from $z_{\text{cond}}^{m-1}$, then,
$z_{\text{cond}}^{m}=\text{Attention}({Q^{\prime}}_{m},{K^{\prime}}_{m},{V^{\prime}}_{m})W^{m}_{\text{O}}.$ (5)
@@
To fuse information of $Z_{\text{cond}}$ into $Z_{\text{music}}$, we compute the cross attention between $Z_{\text{cond}}$ and $Z_{\text{music}}$,
${\rm CrossAttention}(Q,K^{\prime},V^{\prime})=\text{Softmax}\left(\frac{(Q+Q^{\prime})(K^{\prime})^{\mathsf{T}}}{\sqrt{d}}\right)V^{\prime}.$ (6)
Finally, the attention output of $Z_{\text{music}}$ is updated as follows,
$O^{\prime\prime}={\rm CrossAttention}(Q,K^{\prime},V^{\prime})W_{\rm O},$ (7)
$O_{\text{fuse}}=O+g\cdot O^{\prime\prime},$ (8)
where $g\in\mathbb{R}$ is a zero-initialized learnable gating factor.
@@
Thus, the total trainable parameters in Instruct-MusicGen include the input embedding $z^{0}_{\rm cond}$, linear layers ${\rm Linear}_{\rm cond}$, learnable position embeddings $\{e_{1},e_{2},....e_{M}\}$, and learnable gating factors $g$.

### v3

@@ Abstract
The task of text-to-music editing, which employs text queries to modify music (e.g. by changing its style or adjusting instrumental components), presents unique challenges and opportunities for AI-assisted music creation. Previous approaches in this domain have been constrained by the necessity to train specific editing models from scratch, which is both resource-intensive and inefficient; other research uses large language models to predict edited music, resulting in imprecise audio reconstruction. In this paper, we introduce Instruct-MusicGen, a novel approach that finetunes a pretrained MusicGen model to efficiently follow editing instructions such as adding, removing, or separating stems. Our approach involves a modification of the original MusicGen architecture by incorporating a text fusion module and an audio fusion module, which allow the model to process instruction texts and audio input concurrently and yield the desired edited music. Remarkably, although Instruct-MusicGen only introduces $\sim$8new parameters to the original MusicGen model and only trains for 5K steps, it achieves superior performance across all tasks compared to existing baselines. This advancement not only enhances the efficiency of text-to-music editing but also broadens the applicability of music language models in dynamic music production environments.  [^1] Code, model weights and demo are available at: https://github.com/ldzhangyx/instruct-musicgen.  [^2] This work was done during Yixiao Zhang’s internship at Sony AI.
@@ 1 Introduction
The rapid advances in text-to-music generation have opened up new possibilities for AI-assisted music creation [1, 2, 3, 4, 5]. This paradigm shift has also sparked a growing interest in developing models that offer greater controllability [6, 7, 8, 9] and editability [10, 11, 12] over the music generation process. In music production, a stem—a mixed group of tracks often related by instrument type (like drums or lead vocals)—is essential for mixing and mastering because it allows producers to isolate, adjust, and manipulate individual elements of a song. Following the definition in MusicMagus [11], “text-to-music editing" involves using textual queries to modify various aspects of a music recording, which can be categorised into two main types: intra-stem editing, which focuses on modifying a single stem (e.g., changing the instrument, timbre, or performance style), and inter-stem editing, which involves altering the relationships among stems (e.g., adding, removing, or separating stems). Our work mainly focuses on the problem of inter-stem editing.
@@
Previous attempts to develop text-based music editing models have encountered several challenges. Some approaches [13, 10] have focused on training specialised editing models from scratch, which is resource-intensive and may not yield results comparable to state-of-the-art music generation models. Other work [14, 12, 15] has sought to leverage existing large language models (LLMs) and MusicGen [2], allowing the LLM to interpret editing instructions without further training the music model. Although this approach offers flexibility, it often lacks the ability to precisely reconstruct the conditional audio, leading to unreliable results. To address these limitations, an ideal solution should harness the knowledge embedded in pretrained models to ensure high-quality audio output while adapting the architecture to accommodate the specific requirements of music editing tasks.
In this paper, we introduce Instruct-MusicGen, a novel approach that applies an instruction-following tuning strategy to the pretrained MusicGen model, enhancing its ability to follow editing instructions effectively without finetuning all its parameters. As shown in Figure 1, by incorporating an audio fusion module based on LLaMA-Adapter [16, 6] and a text fusion module based on LoRA [17] into the original MusicGen architecture, we allow the model to process both precise audio conditions and text-based instructions simultaneously, which the original MusicGen does not do. This enables Instruct-MusicGen to perform a range of editing tasks. In this paper, we focus on a specific set of these tasks: adding, separating, and removing stems. To train Instruct-MusicGen, we synthesize an instructional dataset using the Slakh2100 dataset [18], introducing only  8% additional parameters compared to the original model, and finetune the model for only 5K steps, which is less than 1% of training a music editing model from scratch.
We evaluate Instruct-MusicGen on two datasets: the Slakh test set and the out-of-domain MoisesDB dataset [19]. Our model outperforms existing baselines and achieves performance comparable to models specifically trained for individual tasks. This demonstrates the effectiveness of our approach in leveraging pretrained models for text-to-music editing while maintaining high-quality results.
@@ 2 Related work
Text-based music editing provides a flexible approach for editing music using textual queries. This method is similar to those used in other modalities that require editing, such as image [20, 21] and video [22, 23] editing. In text-to-music editing, text is used to specify precise alterations to existing music compositions. Previous research such as AUDIT [13] and InstructME [10] developed a diffusion model trained with paired music editing data. Additionally, models like M2UGen [12], Loop Copilot [14], MusicAgent [24], ComposerX [25] and WavCraft [26] use large language models (LLMs) for reasoning and regenerate music with external music generation models. Furthermore, GMSDI [27] attempts to model a joint multi-stem distribution of music for text-based generation and separation. Certain models focus exclusively on specific tasks within music editing, such as conditioned generation [6, 9, 7] and separation [28], along with intra-stem editing tasks such as text-based timbre transfer and style transfer [11, 29, 30, 31].
The task of inter-stem music editing is closely related to stem-wise music generation. Although not directly tied to text-based controls, some research focuses on modeling stem-wise representations to enable simultaneous stem generation and separation. For instance, Jen-1 Composer [32] and MSDM [33] jointly model the distribution of music with four stems using a diffusion model. The abilities of most existing stem-wise music models are restricted to a fixed set of 4 stems, which limits flexibility but enhances controllability. Besides, StemGen [34] trains a LLaMA-based auto-regressive model for flexible stem-wise audio generation.
Our work distinguishes itself from these existing efforts in several key ways. First, rather than developing a new model from scratch or strictly adhering to a fixed set of stems, we leverage the power of a pretrained music language model, MusicGen, and enhance it with instruction tuning. This approach not only reduces the computational cost but also retains the high audio quality of the original MusicGen model. Furthermore, our method introduces minimal additional parameters and requires significantly less training, demonstrating a more efficient and scalable solution for text-based music editing.
@@ 3.1 MusicGen
The original MusicGen consists of three components: (1) the EnCodec [35] audio encoder and decoder, which compress music audio waveforms into latent codes and reconstruct them back into waveforms; (2) a multi-layer transformer architecture that models sequences of latent codes, capturing higher-level music representations and efficiently modeling internal relationships within music audio; and (3) the T5 [36] text encoder, which converts text descriptions into embeddings for text-conditioned generation.
EnCodec employs Residual Vector Quantization (RVQ) [37] to compress audio into tokens using multiple codebooks, where each quantizer encodes the quantization error from the previous one. For a reference audio $X\in\mathbb{R}^{d\cdot f_{n}}$, where $d$ is the duration and $f_{n}$ is the sample rate, EnCodec compresses $X$ into $Q\in\{1,...,L\}^{N\times d\cdot f_{s}}$, where $L$ is the RVQ codebook size, $N$ is the number of codebooks, and $f_{s}$ is the latent code sample rate ($f_{s}<<f_{n}$). In MusicGen, $N=4$, $f_{n}=50$, $f_{s}=32000$, and $L=2048$. Finally, the transformer models the sequence relationships over latent codes [^3] MusicGen’s encodec uses a 50Hz sample rate, which is different from the original 75Hz EnCodec model..
@@ 3.2 Instruct-MusicGen
Instruct-MusicGen takes a music audio input $X^{\text{cond}}$ and a text instruction $X^{\text{instruct}}$ (e.g., "Add guitar") as inputs. The model then edits the music audio $X^{\text{cond}}$ according to the instruction $X^{\text{instruct}}$ and generates the desired edited music $X^{\text{music}}$. As illustrated in Figure 2, Instruct-MusicGen incorporates two additional modules into the vanilla MusicGen: an audio fusion and a text fusion module.
Figure 2: Illustration of the fusion mechanism inside the Transformer module of instruct-MusicGen. The audio fusion module transforms the conditional music audio into embeddings using a duplicated encoder and integrates these embeddings into the MusicGen decoder. The text fusion module modifies the cross-attention mechanism to handle text instructions by finetuning specific layers while keeping the text encoder parameters frozen.
@@ 3.2.1 Audio Fusion Module
The audio fusion module enables Instruct-MusicGen to accept external audio inputs, which is inspired by LLaMA-Adapter [16] and Coco-mulla [6]. The lower part of Figure 2 illustrates the audio fusion module. Initially, we convert $X^{\text{cond}}$ into EnCodec tokens, followed by re-encoding these tokens into the embedding $z^{\text{cond}}$ through the pre-trained embedding layers of MusicGen. Similarly, we transform $X^{\text{music}}$ into the pretrained embedding $z^{\text{music}}$.
@@
The module begins with duplicating self-attention modules of the pretrained MusicGen model to extract latent representations of $z^{\text{cond}}$. Given that MusicGen consists of $M$ layers, we denote
$Z^{\text{cond}}=\{z^{\text{cond}}_{0},z^{\text{cond}}_{1},\ldots,z^{\text{cond}}_{M}\},$ (1)
$Z^{\text{music}}=\{z^{\text{music}}_{0},z^{\text{music}}_{1},\ldots,z^{\text{music}}_{M}\},$ (2)
which represent the hidden states of $X^{\text{cond}}$ and $X^{\text{music}}$ respectively. Note that we use a learnable input embedding as $z^{\text{cond}}_{0}$ and initialize $z^{\text{music}}_{0}$ with $z^{\text{music}}$.
We compute the vanilla self attention for $X^{\text{music}}$ as follows:
$Q^{\text{music}}_{l},K^{\text{music}}_{l},V^{\text{music}}_{l}=\text{QKV-projector}(z^{\text{music}}_{l}),$ (3)
$o^{\text{music}}_{l}=\text{SelfAttn}(Q^{\text{music}}_{l},K^{\text{music}}_{l},V^{\text{music}}_{l}).$ (4)
We project $z^{\text{cond}}$ to a high-dimension representation $h$ via a linear layer $f_{l}$ and learnable positional encoding $e_{l}$,
$h=f_{l}(z^{\text{cond}})+e_{l}.$ (5)
Then, we compute the $(l+1)$-th layer hidden states of $X^{\text{cond}}$ as follows:
$Q^{\text{cond}}_{l},K^{\text{cond}}_{l},V^{\text{cond}}_{l}=\text{QKV-projector}(z^{\text{cond}}_{l}+h),$ (6)
$z^{\text{cond}}_{l+1}={\rm SelfAttn}(Q^{\text{cond}}_{l},K^{\text{cond}}_{l},V^{\text{cond}}_{l}).$ (7)
@@
To fuse information of $X^{\text{cond}}$ into $X^{\text{music}}$, we compute the cross attention between them,
$s^{\text{music}}_{l}=\text{CrossAttn}(Q^{\text{music}}_{l}+Q^{\text{cond}}_{l},K^{\text{cond}}_{l},V^{\text{cond}}_{l}).$ (8)
Finally, the attention output of $X^{\text{music}}$ is updated as follows,
$s^{\prime}_{l}=o^{\text{music}}_{l}+g_{l}\cdot s^{\text{music}}_{l},$ (9)
$z^{\text{music}}_{l+1}=\text{TextFusion}(s^{\prime}_{l},X^{\text{instruct}}),$ (10)
where $g$ is a zero-initialized learnable gating factor.
@@
Thus, the total trainable parameters in Instruct-MusicGen include the input embedding $z^{\rm cond}_{0}$, linear layers $f_{l}$, learnable position embeddings $e_{l}$, learnable gating factors $g$, and learnable parameters in the text fusion module.